OpenStudy (anonymous):

Solve the system algebraically. 2x+ y - 10 = 0 x - y - 4 = 0 What is the value of y? 1/3 2/3 14/3

OpenStudy (anonymous):

will give medal

OpenStudy (solomonzelman):

I add the systems. 2x + y - 10 = 0 + x - y - 4 = 0 ------------------ 3x + 0y -14 = 0 or same as, 3x-14=0 Solve for x, and then solve for y.

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

2 / 3